OpenStudy (anonymous):

Let f(x) = x2 + 6 and g(x) =x+8/x. Find ( g o f)(-7)

OpenStudy (solomonzelman):

can you find the gof ? gof is the other way though, you plug in what the F(x) is equal to, for x, into the G(x). When you find Gof, plug in -7 for x, and there will be (gof)(-7).

OpenStudy (solomonzelman):

If something doesn't make sense to you, please ask.

OpenStudy (anonymous):

i don't undertstand any of it i'm homeschooled

OpenStudy (solomonzelman):

okay, lets just look at each of the functions separately. Okay?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

So lets first see the f(x). \(\large\color{midnightblue}{ f(x)=x^2+6 }\) and in this case, when you want to find f(1) you plug in 1 for x. \(\large\color{midnightblue}{ f(x)=x^2+6 }\) \(\large\color{midnightblue}{ f(1)=1^2+6 }\) \(\large\color{midnightblue}{ ~~~~~~~=1+6 }\) \(\large\color{midnightblue}{ ~~~~~~~=7 }\) So, \(\large\color{midnightblue}{ f(1)=7 }\)

OpenStudy (solomonzelman):

Can you find f(3) ?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

can you plug in 3 for x ?

OpenStudy (anonymous):

how do i do that... i'm sorry i'm lost

OpenStudy (solomonzelman):

We know that \(\large\color{midnightblue}{ f(x)=x^2+6 }\) but look what is happening what I try to do f(3) I replace the x, with the 3. \(\large\color{midnightblue}{ f(3)=3^2+6 }\) good so far?

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

can you tell me what f(3) is going to be equal to ?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

Do you know what it means to say \(\huge\color{midnightblue}{ 3^2 }\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3 times 2?

OpenStudy (solomonzelman):

no

OpenStudy (anonymous):

thats how i remember it... thats what my old teacher said

OpenStudy (solomonzelman):

3 times 2 would be \(\large\color{midnightblue}{3 \times 2 }\) \(\large\color{midnightblue}{3(2) }\) \(\large\color{midnightblue}{(3)2 }\) \(\large\color{midnightblue}{(3)(2) }\) HOWEVER, \(\large\color{midnightblue}{3^2=3 \times 3=9}\)

OpenStudy (anonymous):

wouldmt that be -9

OpenStudy (solomonzelman):

when you say number with a two on top in-front of it, such as \(\large\color{midnightblue}{3^2=3 \times 3~~~~~~~~~~~~~~=9}\) or other examples, \(\large\color{midnightblue}{4^2=4 \times 4~~~~~~~~~~~~~~=16}\) \(\large\color{midnightblue}{1^2=1 \times 1~~~~~~~~~~~~~~=1}\) see?

OpenStudy (solomonzelman):

Just 9, not -9. There is no reason for it to be negative, since all you are doing is that you are multiplying 2 positive numbers together. 3 and 3.

OpenStudy (anonymous):

oh ok i get the examples now

OpenStudy (anonymous):

i just needed the help because i'm on a timer & i couldnt figure it out

OpenStudy (solomonzelman):

So f(3) equals ? (when \(\large\color{midnightblue}{f(x)=x^2+6 }\) )

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

tell me (please) what f(3) will be

OpenStudy (anonymous):

9

OpenStudy (anonymous):

sorry 15?

OpenStudy (solomonzelman):

yes 15.

OpenStudy (solomonzelman):

Can you find f(-2) ?

OpenStudy (anonymous):

no